# Investigating Disordered Granular Matter via Ordered Geometric Fragmentation

### v1

@@ Abstract
The evolution of occupied volume under progressive fragmentation of granular matter is studied using a purely geometric model. Rather than modeling disorder directly, properties of disordered granular assemblies are investigated by analyzing an associated family of highly ordered reference configurations that provide sharp upper bounds on accessible volume. Grains are idealized as fragments derived from a hypothetical elongated parent prism with square cross section, sequentially sliced and reassembled into configurations that maximize the enclosed volume.
Analytic expressions are derived for the maximal attainable volume at each fragmentation stage. The volume evolution is non-monotonic: initial fragmentation produces structures whose volume exceeds that of the original object, while further fragmentation leads to monotonic decrease converging to a limiting value of 5/4 times the initial volume, independent of the total number of fragments.
The model reveals pairs of reassembled configurations built from geometrically indistinguishable building blocks yet enclosing different volumes. These conjugate configurations constitute purely geometric analogues of distinct phases connected by rearrangement-induced transitions. Explicit relations link the idealized construction to experimentally measurable grain parameters. Comparison with experimental data on cylindrical rods shows the predicted upper bound falls within the observed packing density range, demonstrating that the model captures essential geometric features despite its simplicity.
@@ I Introduction
In contrast, the present work addresses a complementary question: how does the volume occupied by granular matter evolve under progressive fragmentation? Rather than modelling disorder directly, this question is examined through a controlled geometric construction based on a single elongated parental object that is sequentially partitioned into identical building blocks. At each fragmentation stage, the maximal enclosed volume attainable after reassembly is evaluated. The resulting configurations are not intended to represent typical granular packings; instead, they provide sharp geometric bounds that elucidate systematic trends in volume evolution induced by fragmentation.
@@
While much attention has been devoted to the packing of particles with fixed shapes, the influence of fragmentation on packing density remains less systematically explored. Fragmentation processes — whether mechanical comminution, crushing, or controlled slicing—alter both the size distribution and the geometric properties of constituent grains [11]. The resulting changes in particle aspect ratio and surface morphology can significantly modify packing efficiency. The present work isolates the purely geometric contribution to these effects by considering idealized fragmentation of a single parental object into progressively smaller, identical building blocks.
Building on these ideas, a minimal, fully analytic model is constructed that isolates the geometric consequences of fragmentation and reassembly. Explicit expressions for the maximal enclosed volume at each fragmentation stage are obtained, and several nontrivial structural features are revealed. In particular, the emergence of paired (conjugate) configurations composed of geometrically indistinguishable building blocks, yet enclosing different volumes, is identified — a phenomenon that has no direct analogue in conventional packing problems.
The objectives of this paper are as follows: (i) to quantify the evolution of volume under sequential fragmentation of granular matter using a simple geometric model; (ii) to investigate structural relations among configurations reassembled from differently fragmented building blocks; and (iii) to derive explicit formulae that connect the parameters of the model to experimentally measurable quantities.
@@ II Model and notation
In this work, the geometric properties of granular matter composed of idealised grains are investigated. Friction between grains is assumed to be negligible, and the weight of individual grains is ignored. All grains are assumed to originate from a single, very long hypothetical parental object. The number of grains produced and their individual sizes are determined entirely by the geometry of this initial object.
Grains are modeled as rectangular parallelepipeds with square cross-section. Accordingly, a family of long rectangular prisms (hereafter referred to as initial objects) is considered, denoted by $T_{\,n}^{\,0}$, with square cross-section $a\times a$ and length
@@
$R_{\,n}^{\,i}=\frac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}.$
@@ II.1 Fragmentation rule
The fragmentation rule is defined recursively as follows:
@@
    Stage 2: Each piece obtained at Stage 1 is divided lengthwise into two equal parts. The resulting eight pieces are reassembled into the tower $T_{\,n}^{\,2}$.
@@
    This procedure is continued by dividing each piece lengthwise by a factor of two until stage $n+1$, at which point all pieces are unit cubes and the maximal-volume configuration $T_{\,n}^{\,n+1}$ is obtained.
@@
The square geometry therefore represents an extremal, volume-maximizing configuration within this class and serves as a natural reference case. No claim is made that this construction exhausts all possible configurations, nor that a global maximal-volume theorem is established.
@@ III Fragmentation – reassembly process and volume evolution
Using the fragmentation rule introduced in the previous section, the slicing of long objects is explicitly presented for the first few values of $n$. For each $n$, the sequence of towers $T_{\,n}^{\,i}$ produced at fragmentation stages $i=0,1,\dots,n+1$ is listed together with their volumes $V_{\,n}^{\,i}$ and corresponding relative volumes $R_{\,n}^{\,i}$. These examples motivate the general formulae, which are subsequently derived using mathematical induction.
In the figures below, thick lines indicate grain boundaries, while thin lines serve only to illustrate the relative sizes of the grains.
@@ III.1 Case $n=0$ ($l=4a$)
In the following, the findings are summarised and rewritten in a form that is convenient for deriving general formulae for the volumes, relative volumes, and the total number of fragmentation stages.
@@
$T(n)$ denotes the total number of fragmentation stages for an initial object indexed by $n$.
@@ III.2 Case $n=1$ ($l=8a$)
$V_{\,1}^{\,2}$ $=8a^{3}+2a^{3}=4\times 2^{1}a^{3}+2^{2\times 1-1}a^{3},$ (4)
$T(1)$ $=1+1.$ (5)
The tower $T_{\,1}^{\,1}$ is built from blocks of length $2a$, whereas the tower $T_{\,1}^{\,2}$ is constructed from unit blocks. These unit blocks are arranged such that the resulting structure appears as if it were built from the same type of building blocks as $T_{\,1}^{\,1}$.
This observation suggests that the tower $T_{\,1}^{\,2}$ may be regarded, at a coarse-grained geometric level, as being assembled from the same building blocks as $T_{\,1}^{\,1}$, differing only in their global arrangement. In this sense, one configuration may be transformed into the other by rearrangement of the building blocks, despite differences in their internal composition.
@@
If no energetic preference is associated with either configuration, both may occur with comparable probability despite their different volumes. In this sense, the two configurations may be viewed as distinct “geometric phases,” one denser than the other.
@@ III.3 Case $n=2$ ($l=16a$)
It is observed that, relative to the initial object, the volume of the towers first increases at the initial fragmentation stage and then decreases monotonically at subsequent stages. This behaviour is governed by the cavity formed during the first rearrangement: while a cavity is created at the first stage, its volume decreases as the fragmentation proceeds.
A further parallel with the previous case may be drawn. The towers $T_{\,2}^{\,1}$ and $T_{\,2}^{\,3}$ may be regarded as being built from identical building elements of length $4a$, arranged in different geometric configurations. As before, such a pair is referred to as a pair of conjugate towers. The intermediate tower $T_{\,2}^{\,2}$ therefore lacks an obvious conjugate partner.
@@
The tower $T_{\,2}^{\,2}$ is therefore referred to as a self-conjugate tower. It will be shown later that, whenever $n$ is even, the middle tower in the sequence is self-conjugate.
From this perspective, transitions between conjugate towers may be interpreted as transitions between distinct “geometric phases” with different densities. In contrast, the transition associated with a self-conjugate tower corresponds to a change between different symmetry realisations of the same geometric state. In this sense, transitions between conjugate towers resemble first-order–like phase transitions, while transitions involving self-conjugate towers may be viewed as second-order–like phase transitions.
@@
$(T_{\,n}^{\,1},T_{\,n}^{\,n+1}),\,(T_{\,n}^{\,2},T_{\,n}^{\,n}),\,(T_{\,n}^{\,3},T_{\,n}^{\,n-1})\dots$
If $n$ is odd, there are $(n+1)/2$ such pairs, and every tower has a conjugate partner. If $n$ is even, there are $n/2$ conjugate pairs, while a single “middle” tower, $T_{n}^{n/2+1}$, remains self-conjugate. This odd – even distinction plays a central role in the classification of geometrical phase transitions and underlies the general formulae derived below.
It is now shown that the lateral faces of a self-conjugate tower are necessarily square. This requires the width and height of a lateral face to be equal. Since the width of a tower face is equal to the length of its building blocks, it is sufficient to demonstrate that this length coincides with the height of the tower.
@@
Substitution of $i=n/2+1$ yields $l_{i}=2^{\,n/2}a$.
Third, the height of the tower is determined by the number of levels multiplied by $a$. A single level is formed at the first stage, and the number of levels doubles at each subsequent stage. The number of levels at stage $i$ is thus $2^{\,i-1}$, therefore height of the tower $h_{i}$ at stage $i$ is
@@
Substitution of $i=n/2+1$ gives a tower height of $2^{\,n/2}a$.
Hence, the width and height of the lateral faces are therefore equal, and the lateral faces of a self-conjugate tower are square. ∎
@@
The number of towers $T(n)$ created from an initial parental object with dimensions $2^{\,n+2}a\times a\times a$ is given by
@@
It then follows directly that the final stage index is $f=n+1$. Consequently, the total number of towers produced during the fragmentation process is equal to $n+1$, which completes the proof. ∎
For each $n$, the sequence of towers $T_{\,n}^{\,i}$ exhibits a natural pairing: $T_{\,n}^{\,1}$ with $T_{\,n}^{\,n+1}$, $T_{\,n}^{\,2}$ with $T_{\,n}^{\,n}$, and so on. Towers within each pair are constructed from identical looking building blocks, but differ by a global rearrangement that leads to different enclosed cavity volumes.
@@
From this relation, the paired fragmentation stage is obtained as p=ni+2p=n-i+2. Consequently, the conjugate tower pairs are $(T_{\,n}^{\,i},T_{\,n}^{\,n-i+2})$. ∎
@@ III.4 General volume formula
The volume of a tower is given by the sum of the material volume $2^{n+2}a^{3}$ of the initial parental object, which remains unchanged throughout the fragmentation process, and the volume of the enclosed space (cavity). The cavity volume is given by the product of the cross-sectional area of the cavity and the height of the tower, with the latter given by Eq. (7). The cross-sectional area of the cavity is equal to the square of the building-block length $l_{i}$, given by Eq. (6). The enclosed cavity volume is therefore expressed as
@@
The volume $V_{\,n}^{\,i}$ of the tower $T_{\,n}^{\,i}$ reassembled at fragmentation stage $i$ from an initial parental object of dimensions $2^{\,n+2}a\times a\times a$ is given by
@@
The corresponding relative volume is therefore obtained as
$R_{\,n}^{\,i}=\dfrac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}=1+2^{\,n-i-1}.$ (10)
The variation of the relative volume as a function of the fragmentation stage $i$ for several values of $n$ is illustrated in Fig. 4. As established above, the final fragmentation occurs at stage $i=n+1$. The relative volume at the final stage is therefore universal and is given by
@@
The value $5/4$ thus represents a universal upper bound for the relative volume obtained after complete fragmentation of any sufficiently long object with square cross section.
The plots further indicate that the first fragmentation stage produces the largest increase in volume, after which the relative volume decreases monotonically with increasing stage index. This behaviour is governed by the evolution of the central cavity: the initial object encloses no void space, whereas the first fragmentation generates the largest possible cavity. At each subsequent stage, the cross-sectional area of the cavity is reduced by a factor of four, while the height of the tower increases only by a factor of two. Consequently, the enclosed cavity volume is halved at each stage, while the relative volume decreases monotonically but is not halved because of the additive constant.
@@
Figure 4: Relative volume $R_{\,n}^{\,i}$ as a function of fragmentation stage $i$ for several values of $n$. Each curve corresponds to a fixed initial length $l=2^{n+2}a$. The first fragmentation stage yields the maximal relative volume, followed by a monotonic decrease towards the universal limit $R=5/4$ at $i=n+1$.

### v2

@@ Abstract
The evolution of occupied volume under progressive fragmentation of granular matter is studied using a purely geometric model. Rather than modelling disorder directly, properties are investigated by analysing highly ordered reference configurations that provide sharp upper bounds on accessible volume. Grains are idealised as fragments from a hypothetical elongated parent prism with square cross section, sequentially sliced and reassembled into configurations that maximise enclosed volume.
Analytic expressions are derived for the maximal volume at each fragmentation stage. Volume evolution is non-monotonic: initial fragmentation produces structures exceeding the original volume, while further fragmentation leads to monotonic decrease converging to 5/4 times the initial volume, independent of fragment number. The packing fraction obeys the asymptotic scaling law of inverse proportionality to aspect ratio, in agreement with experimental observations.
The model reveals pairs of configurations built from geometrically indistinguishable building blocks yet enclosing different volumes. These conjugate configurations constitute geometric analogues of distinct phases connected by rearrangement-induced transitions. A criterion for observability is derived, showing such transitions are restricted to systems of limited grain number but may occur locally as domain formation in larger assemblies. Comparison with experimental data confirms the model provides a lower bound on packing fraction and predicts domain sizes should scale linearly with aspect ratio, testable through X-ray tomography.
@@ I Introduction
Recent theoretical progress has been made in understanding random close packing (RCP) of monodisperse spherical particles through analytical frameworks linking packing fraction to coordination number and the onset of mechanical rigidity [4]. These approaches identify the jamming transition as a critical point characterized by a threshold coordination number, with packing fractions that are largely independent of system size. However, the case of non-spherical elongated particles remains analytically challenging and has been explicitly identified as an open problem requiring further investigation [4]. The present work contributes to this area through a purely geometric approach that reveals how fragmentation-induced volume evolution in elongated grain assemblies can be understood from geometric constraints alone, without recourse to mechanical or energetic considerations.
In contrast to conventional packing studies, the present work addresses a complementary question: how does the volume occupied by granular matter evolve under progressive fragmentation? Rather than modelling disorder directly, this question is examined through a controlled geometric construction based on a single elongated parent object that is sequentially partitioned into identical building blocks. At each fragmentation stage, the maximal enclosed volume attainable after reassembly is evaluated. The resulting configurations are not intended to represent typical granular packings; instead, they provide sharp geometric bounds that elucidate systematic trends in volume evolution induced by fragmentation.
@@
While much attention has been devoted to the packing of particles with fixed shapes, the influence of fragmentation on packing density remains less systematically explored. Fragmentation processes — whether mechanical comminution, crushing, or controlled slicing — alter both the size distribution and the geometric properties of constituent grains [12]. The resulting changes in particle aspect ratio and surface morphology can significantly modify packing efficiency. The present work isolates the purely geometric contribution to these effects by considering idealized fragmentation of a single parent object into progressively smaller, identical building blocks.
Building on these ideas, a minimal, fully analytic model is constructed that isolates the geometric consequences of fragmentation and reassembly. Explicit expressions for the maximal enclosed volume at each fragmentation stage are obtained, and several structural features are revealed. In particular, the emergence of paired (conjugate) configurations composed of geometrically indistinguishable building blocks, yet enclosing different volumes, is identified. The model predicts bounds on packing density that are shown to be consistent with experimental scaling laws for elongated particles. Furthermore, the geometric framework reveals local ordering phenomena in bulk granular assemblies that can be tested through X-ray tomography experiments.
The objectives of this paper are as follows: (i) to quantify the evolution of volume under sequential fragmentation of granular matter using a simple geometric model; (ii) to investigate structural relations among configurations reassembled from differently fragmented building blocks; (iii) to derive explicit formulae that connect the parameters of the model to experimentally measurable quantities; and (iv) to compare theoretical predictions with experimental data and propose testable signatures of the geometric mechanisms identified here.
@@ II Model and notation
In this work, the geometric properties of granular matter composed of idealised grains are investigated within a purely static geometric framework. Friction between grains is assumed to be negligible, and the weight of individual grains is ignored. All grains are assumed to originate from a single, very long hypothetical parent object. The number of grains produced and their individual sizes are determined entirely by the geometry of this initial object.
Grains are modelled as rectangular parallelepipeds with square cross-section. Accordingly, a family of long rectangular prisms (hereafter referred to as initial objects) is considered, denoted by $T_{\,n}^{\,0}$, with square cross-section $a\times a$ and length
@@
$R_{\,n}^{\,i}=\frac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}.$ (4)
In this notation, the lower index $n$ characterises the size of the initial parent object via Eq. (1), while the upper index $i$ labels the fragmentation stage.
@@ II.1 Fragmentation rule
The fragmentation rule is defined recursively as follows. Stage 0 denotes the initial configuration prior to any fragmentation, while stages $i=1,2,\dots,n+1$ correspond to successive fragmentation steps.
@@
    Stage 2: Each piece obtained at Stage 1 is divided along its longest axis into two equal parts. The resulting eight pieces are reassembled into the tower $T_{\,n}^{\,2}$.
@@
    This procedure is continued by dividing each piece along its longest axis into two equal parts at each subsequent stage, until stage $n+1$, at which point all pieces are unit cubes and the maximal-volume configuration $T_{\,n}^{\,n+1}$ is obtained.
@@
The square geometry therefore represents an extremal, volume-maximising configuration within this class and serves as a natural reference case. No claim is made that this construction exhausts all possible configurations, nor that a global maximal-volume theorem is established.
@@ III Fragmentation – reassembly process and volume evolution
Using the fragmentation rule introduced in Sec. II.1, the slicing of long objects is presented explicitly below for the first few values of $n$. For each $n$, the sequence of towers $T_{\,n}^{\,i}$ produced at fragmentation stages $i=1,\dots,n+1$, including the initial object $i=0$, is listed alongside their volumes $V_{\,n}^{\,i}$ and corresponding relative volumes $R_{\,n}^{\,i}$. These examples motivate the general formulae, which are subsequently derived using mathematical induction.
In the figures that follow, thick lines indicate grain boundaries, while thin lines serve only to illustrate the relative sizes of the grains.
@@ III.1 Case $n=0$ ($l=4a$)
In the following, the findings are summarised and rewritten in a form that is convenient for deriving general formulae for the volumes, relative volumes, and the total number of fragmentation stages. $T(n)$ denotes the total number of fragmentation stages for an initial object indexed by $n$, so that
@@ III.2 Case $n=1$ ($l=8a$)
$V_{\,1}^{\,2}$ $=8a^{3}+2a^{3}=4\times 2^{1}a^{3}+2^{2\times 1-1}a^{3},$
$T(1)$ $=1+1.$
The tower $T_{\,1}^{\,1}$ is built from blocks of length $2a$, whereas the tower $T_{\,1}^{\,2}$ is constructed from unit blocks arranged such that the resulting structure appears, at a coarse-grained geometric level, as if it were assembled from the same building blocks as $T_{\,1}^{\,1}$, differing only in their global arrangement. In this sense, one configuration may be transformed into the other by rearrangement of the building blocks, despite differences in their internal composition.
@@
Within the purely geometric framework of the present model, if no energetic preference is associated with either configuration, both may occur with comparable probability despite their different volumes. In this sense, the two configurations may be viewed as distinct ‘geometric phases’, one denser than the other.
@@ III.3 Case $n=2$ ($l=16a$)
Relative to the initial object, the volume of the towers first increases at the initial fragmentation stage and then decreases monotonically at subsequent stages. This behaviour is governed by the cavity formed during the first rearrangement: while a cavity is created at the first stage, its volume decreases as the fragmentation proceeds.
A further parallel with the case $n=1$ (Sec. III.2) may be drawn. The towers $T_{\,2}^{\,1}$ and $T_{\,2}^{\,3}$ may be regarded as being built from identical building elements of length $4a$, arranged in different geometric configurations. As before, such a pair is referred to as a pair of conjugate towers. The intermediate tower $T_{\,2}^{\,2}$ does not possess a conjugate partner in the same sense.
@@
The tower $T_{\,2}^{\,2}$ is therefore referred to as a self-conjugate tower. As shown in Proposition 1 below, whenever $n$ is even, the middle tower in the sequence is self-conjugate.
From this perspective, transitions between conjugate towers may be interpreted as transitions between distinct ‘geometric phases’ with different densities. In contrast, the transition associated with a self-conjugate tower corresponds to a change between different symmetry realisations of the same geometric state. In this sense, transitions between conjugate towers resemble first-order-like phase transitions, while transitions involving self-conjugate towers may be viewed as second-order-like phase transitions.
@@
$(T_{\,n}^{\,1},T_{\,n}^{\,n+1}),\,(T_{\,n}^{\,2},T_{\,n}^{\,n}),\,(T_{\,n}^{\,3},T_{\,n}^{\,n-1}),\,\dots$ (5)
If $n$ is odd, there are $(n+1)/2$ such pairs, and every tower has a conjugate partner. If $n$ is even, there are $n/2$ conjugate pairs, while a single ‘middle’ tower, $T_{\,n}^{\,n/2+1}$, remains self-conjugate. This odd–even distinction plays a central role in the classification of geometrical phase transitions and underlies the general formulae derived in Sec. III.4.
The following proposition establishes that the lateral faces of a self-conjugate tower are necessarily square, as claimed above. This requires the width and height of a lateral face to be equal. Since the width of a tower face is equal to the length of its building blocks, it is sufficient to demonstrate that this length coincides with the height of the tower.
@@
Substitution of $i=n/2+1$ into Eq. (6) yields $l_{\,n/2+1}=2^{\,n/2}a$.
Third, the height of the tower is determined by the number of levels multiplied by $a$. A single level is formed at the first stage, and the number of levels doubles at each subsequent stage. The number of levels at stage $i$ is thus $2^{\,i-1}$, therefore the height of the tower $h_{i}$ at stage $i$ is
@@
Substitution of $i=n/2+1$ into Eq. (7) gives $h_{\,n/2+1}=2^{\,n/2}a$.
Hence, the width and height of the lateral faces are equal, and the lateral faces of a self-conjugate tower are square. ∎
@@
The number of towers $T(n)$ created from an initial parent object with dimensions $2^{\,n+2}a\times a\times a$ is given by
@@
It then follows directly that the final stage index is $f=n+1$. Since towers are produced at stages $i=1,2,\dots,f$, the total number of towers produced during the fragmentation process is equal to $n+1$, which completes the proof. ∎
For each $n$, the sequence of towers $T_{\,n}^{\,i}$ exhibits a natural pairing: $T_{\,n}^{\,1}$ with $T_{\,n}^{\,n+1}$, $T_{\,n}^{\,2}$ with $T_{\,n}^{\,n}$, and so on. Towers within each pair are constructed from identical-looking building blocks, but differ by a global rearrangement that leads to different enclosed cavity volumes. This pairing is made precise in the following corollary.
@@
From this relation, the paired fragmentation stage is obtained as p=ni+2p=n-i+2. Substituting back, the conjugate tower pairs are $(T_{\,n}^{\,i},T_{\,n}^{\,n-i+2})$, which completes the proof. ∎
@@
Proof.
The result follows directly from Corollary 2 by counting the pairs. ∎
@@ III.4 General volume formula
The volume of a tower is given by the sum of the material volume $2^{n+2}a^{3}$ of the initial parent object, which remains unchanged throughout the fragmentation process, and the volume of the enclosed space (cavity). The cavity volume is given by the product of the cross-sectional area of the cavity and the height of the tower, with the latter given by Eq. (7). The cross-sectional area of the cavity is equal to the square of the building-block length $l_{i}$, given by Eq. (6). The enclosed cavity volume is therefore expressed as
@@
The total tower volume then follows immediately.
@@
The volume $V_{\,n}^{\,i}$ of the tower $T_{\,n}^{\,i}$ reassembled at fragmentation stage $i$ from an initial parent object of dimensions $2^{\,n+2}a\times a\times a$ is given by
@@
The corresponding relative volume is obtained as
$R_{\,n}^{\,i}=\frac{V_{\,n}^{\,i}}{V_{\,n}^{\,0}}=\frac{2^{\,n+2}a^{3}+2^{\,2n-(i-1)}a^{3}}{2^{\,n+2}a^{3}}=1+2^{\,n-i-1}.$ (10)
The variation of the relative volume as a function of the fragmentation stage $i$ for several values of $n$ is illustrated in Fig. 4. As established in Corollary 1, the final fragmentation occurs at stage $i=n+1$. The relative volume at the final stage is therefore independent of $n$ and is given by
@@
The value $5/4$ thus represents a geometric upper bound for the relative volume obtained after complete fragmentation within the present model.
The plots further indicate that the first fragmentation stage produces the largest increase in volume, after which the relative volume decreases monotonically with increasing stage index. This behaviour is governed by the evolution of the central cavity: the initial object encloses no void space, whereas the first fragmentation generates the largest possible cavity. At each subsequent stage, the cross-sectional area of the cavity is reduced by a factor of four, while the height of the tower increases only by a factor of two. Consequently, the enclosed cavity volume — and hence the cavity contribution to the relative volume — is reduced by a factor of two at each stage. The relative volume itself decreases monotonically but does not halve, owing to the additive contribution of the material volume.
@@
Figure 4: Relative volume $R_{\,n}^{\,i}$ as a function of fragmentation stage $i$ for several values of $n$. Each curve corresponds to a fixed initial length $l=2^{n+2}a$. The first fragmentation stage yields the maximal relative volume, followed by a monotonic decrease towards the limiting value $R=5/4$ at $i=n+1$.